OpenStudy (anonymous):

Every female chicken lays 4 eggs. What is the constant of proportionality for the ratio of eggs to female chickens? 4 3 one-third one-fourth

OpenStudy (anonymous):

@Brostep0s @bears55

OpenStudy (anonymous):

@Brostep0s @shinalcantara

OpenStudy (anonymous):

so then i think it would be 1/4 right?

OpenStudy (anonymous):

ok, thanks!

OpenStudy (shinalcantara):

since it's eggs to female then 4/1 = 4

OpenStudy (anonymous):

so then the answer would be 4?

OpenStudy (shinalcantara):

always take note on which comes first in the statement

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@shinalcantara is that right?

OpenStudy (anonymous):

@shinalcantara?

OpenStudy (anonymous):

so the answer is either 1/4 or 4. which one?

OpenStudy (anonymous):

@Brostep0s what do you think?

OpenStudy (shinalcantara):

lemme recheck first.

OpenStudy (anonymous):

I would think that since it says that one chicken lays four eggs, that it would be compared as a whole, whereas 4:1 would be the ratio.

OpenStudy (anonymous):

It doesn't give any certain amount of chickens, so there is not any simplifying to do.

OpenStudy (anonymous):

but that's not in the answer choices. the answers are: 4 3 one-third one-fourth

OpenStudy (anonymous):

i think the ratio is 1:4 but i dont know what the constant of proportionality is

OpenStudy (shinalcantara):

for direct proportion since it's eggs to females, then the independent variable x is the female and y which is the eggs is the dependent variable y = kx 4 = k (1) k = 4

OpenStudy (anonymous):

its just 4 @77777jeannie77777

OpenStudy (anonymous):

oh ok thanks! can you guys help me with one more?

OpenStudy (anonymous):

that's what 4:1 means, it comes in different forms, "4:1, 4/1, 4, 4.0, Etc."

OpenStudy (anonymous):

& yes

OpenStudy (anonymous):

The graph shows the amount of money paid when purchasing bags of caramel corn at the zoo: What is the constant of proportionality? 3 one-third 2 one-half

OpenStudy (anonymous):

i think i have a 50/50 chance because i think its either 3 or 2

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

ok, can you help with one more?

OpenStudy (shinalcantara):

the constant of proportionality is just the slope of the line.. y/x = 3/1 = 3

OpenStudy (anonymous):

The graph shows the amount of money paid when purchasing bags of peanuts at the zoo: Write an equation to represent the relationship between the total cost (y) and the number of bags of peanuts (x). y = 4x y = one-fourth x y = 3x y = one-third x

OpenStudy (anonymous):

yes, Slope, or "rate of change" can be found by, in this case, being compared to \[Rise/Run\]

OpenStudy (anonymous):

y=4x

OpenStudy (anonymous):

thanks! one more pls

OpenStudy (anonymous):

last one :D

OpenStudy (anonymous):

Ellen bought gifts for her friends, and each gift cost $7. Let y represent the total cost and x represent the number of gifts. Write an equation to represent the proportional relationship between the total cost and the number of gifts she bought. y = 7x y = one-seventh x y = 6x y = one-sixth x

OpenStudy (anonymous):

y = 7x right?

OpenStudy (anonymous):

y=7x. It's a simple plug-in: match-up formula.

OpenStudy (anonymous):

thanks! medal and fan have been given

OpenStudy (anonymous):

bye!

OpenStudy (anonymous):

yeppers cx